# Never-Ending Embodied Robot Learning

### v1

@@ 3.2 Skill-Specific Evolving Planner
0:  Robotics incremental tasks $\{\mathcal{T}^{m}\}_{m=1}^{M}$ with datasets $\mathcal{T}^{m}=\{\mathcal{D}^{m}\}_{i=1}^{N^{d}}$; Initialized perceiver model: $\Theta^{0}_{p}$; Initialized NeRF model: $\Theta^{0}_{n}$; Pre-trained diffusion model: $\Theta_{u}$; Pre-trained CLIP language encoder: $\mathcal{E}_{c}$; Initialized memory buff: $\mathcal{M}=\emptyset$; Iterations: $\{\mathcal{I}^{m}\}_{m=1}^{M}$.
@@
5:     Compute $\mathcal{L}_{ssr}$ by SSR ($\mathrm{v}^{m},\mathrm{v}^{m-1},\mathrm{l}_{x},\Theta_{u}$) using Eq. (4) and Eq. (12);
6:     $\mathbf{S}[I,:],\mathbf{W}[I,:]\xleftarrow{}$ ESP ($\mathbf{l}_{s}$);
7:     Compute $\mathcal{L}_{ce},\mathcal{L}_{srd}$ utilizing $\mathbf{S}[I,:]$, $\mathbf{W}[I,:]$, $\Theta^{m}_{p},\Theta^{m-1}_{p}$;
@@
3.3 Skill-Shared Semantics Rendering Module
For language-conditional behaviour-cloning manipulation, a comprehensive semantics understanding of the 3D scene (Driess et al., 2022) plays a key role in enabling agent to perform complicated manipulation skills. Especially in NERL problem, there exist skill-shared semantics across various skills, such as 3D object and scene semantics. The existence of forgetting on semantic space makes these semantics incomplete, further resulting in catastrophic forgetting on old skills. Considering this motivation, we develop a skill-shared semantics rendering module (SSR) to transfer skill-shared semantic information of 3D voxel space, where a NeRF model and vision foundation model provide semantics supervision to effectively enrich the 3D voxel semantics. Drawing inspiration from 3D visual representation learning methods (Shim et al., 2023; Ze et al., 2023), we leverage a latent-conditioned NeRF architecture (Yu et al., 2021) not only to synthesizes RGB color $\mathbf{c}$ of a novel image views like traditional NeRF (Mildenhall et al., 2021), but also to render the semantic feature $\mathbf{s}$ from 3D voxel space as follows:
where $\mathcal{F}_{\Theta^{m}_{n}}$ denotes the neural rendering function of NeRF model $\Theta^{m}_{n}$. The 3D voxel feature $\mathbf{v}_{s}$ is obtained by a grid sample method based on trilinear interpolation from the 3D voxel observation $\mathbf{v}$. $\mathbf{x}$, $\sigma$ is the 3D input point and differential density, and $\mathbf{d}$ represents unit viewing direction. The camera ray $\mathbf{r}$ can be obtained by: $\mathbf{r}=\mathbf{o}+t\mathbf{d}$, where $\mathbf{o}$ indicates the camera origin. By adding field-wise branches, $\Theta^{m}_{n}$ performs the same neural rendering function $\mathcal{F}_{\Theta^{m}_{n}}$ to estimate RGB color $\mathbf{c}$ and semantic feature $\mathbf{s}$. Thus, the same accumulated transmittance $T(t)$ is shared to predict the two different fields and is defined as follows:
[…]
In light of this, a RGB image $\mathbf{C}$ and 2D semantic map $\mathbf{S}$ can be rendered as :
[…]
To distill skill-shared knowledge, we initialize a NeRF model acquired from the last task and denote it as $\Theta_{n}^{m-1}$. Then, we feed the same input to obtain the pseudo ground truth $\hat{\mathbf{C}}$ by Eq. 6. We design a loss function to supervise the reconstruction process as follows:
$\mathcal{L}_{c}$ $=\underset{\mathbf{r}\in\mathcal{R}}{\sum}\|\mathbf{C}(\mathbf{r},\mathbf{v}_{s})-\mathbf{Y}_{c}(\mathbf{r})\|^{2}_{2}$
$+\beta\underset{\mathbf{r}\in\mathcal{R}}{\sum}\|\mathbf{C}(\mathbf{r},\mathbf{v}_{s})-\hat{\mathbf{C}}(\mathbf{r},\mathbf{v}_{s})\|^{2}_{2}\cdot\mathbb{I}_{\mathbf{v}_{s}\notin\mathcal{T}^{m}},$ (8)
where $\mathbf{Y}_{c}$ indicates the ground truth color and $\mathcal{R}$ is the set of all camera rays. $\alpha$ is the hyper-parameter to control the weight of loss function. $\mathbb{I}_{\mathbf{v}_{s}\notin\mathcal{T}^{m}}$ is defined such that $\mathbb{I}_{\mathbf{v}_{s}\notin\mathcal{T}^{m}}=1$ when the condition $\mathbf{v}_{s}\notin\mathcal{T}^{m}$ is satisfied, and $\mathbb{I}_{\mathbf{v}_{s}\notin\mathcal{T}^{m}}=0$ otherwise.
@@ 3.3 Skill-Shared Semantics Rendering Module
Figure 3: Visualization of prediction results on various manipulation skills between Ours, Ours-w/oSRD and Ours-w/oSEP&SRD.
Considering the insufficiency in capturing skill-shared semantics by reconstructing novel views, we introduce a pre-trained visual foundation model that contains robust scene semantics to provide supervision. Relying on being pre-trained on large-scale vision-language dataset, Stable Diffusion model (Rombach et al., 2022) can possess robust intrinsic representational capabilities and are consequently utilized for semantic representation in segmentation and classification tasks (Xu et al., 2023; Li et al., 2023). In light of this, we employ a text-to-image Stable Diffusion model $\Theta_{u}$ to extract vision-language semantic feature for supervising. Specifically, given a input view, i.e., $\mathbf{Y}_{c}$, we perform a one-step noise adding process to obtain a a noisy image $\mathbf{Y}_{c,t}$. Then we utilize our diffusion model $\Theta_{u}$ to collect vision-language semantic feature as the ground truth $\hat{\mathbf{S}}$:
$\mathbf{Y}_{c,t}(\mathbf{r}):=\sqrt{\alpha_{t}}\mathcal{E}_{v}(\mathbf{Y}_{c}(\mathbf{r}))+\sqrt{1-\alpha_{t}}\epsilon,$ (9)
$\hat{\mathbf{S}}(\mathbf{r},\mathbf{l}_{p})=\Theta_{u}(\mathbf{Y}_{c,t}(\mathbf{r}),\mathcal{E}_{c}(\mathbf{l}_{p})),$ (10)
where $\mathcal{E}_{v}$ is a VAE encoder to encode image $\mathbf{Y}_{c}$ from pixel space to latent semantic space. $t$ represents the diffusion process step, $\epsilon\sim\mathcal{N}(0,1)$ and $\alpha_{t}$ is designed to control the noise schedule. $\mathbf{l}_{p}$ denotes the language prompt modified from task description $\mathbf{l}$. To this end, we align the rendered semantic feature $\mathbf{S}$ and diffusion feature $\hat{\mathbf{S}}$ to perform semantics transfer as follows:
$\mathcal{L}_{s}=\underset{\mathbf{r}\in\mathcal{R}}{\sum}\|\mathbf{S}(\mathbf{r},\mathbf{v}_{s})-\hat{\mathbf{S}}(\mathbf{r},\mathbf{l}_{p})\|^{2}_{2}.$ (11)
In summary, the major objective of our SSR module to complete skill-shared semantics can be expressed as:
$\mathcal{L}_{srr}=\mathcal{L}_{c}+\lambda_{1}\mathcal{L}_{s},$ (12)
where $\lambda_{1}$ is the hyper-parameter.
@@
Table 1: Comparisons of success rate (%) on Kitchen and Living Room. Red and Blue represents the highest results and runner-up.
@@
3.4 Skill-Shared Representation Distillation Module
To address catastrophic forgetting on old skills, we develop a skill-shared representation distillation module (SRD) to align skill-shared representation, as presented in Fig. 2. Specifically, given a multi-modal keyframe input $\mathbf{k}^{m}_{j}=\{\mathbf{a}^{m}_{j},\mathbf{r}^{m}_{j},\mathbf{l}^{m}\}$ from a mini batch, we can obtain a keyframe prediction $\mathbf{P}^{m}(\mathbf{k}^{m}_{j},\Theta^{m}_{p})=\{\mathbf{P}^{m}_{j,tran},\mathbf{P}^{m}_{j,rot},\mathbf{P}^{m}_{j,grip},\mathbf{P}^{m}_{j,col}\}$. To supervise our NERagent to learn skill-wise knowledge, we follow (Shridhar et al., 2023) to introduce a cross-entropy loss for optimizing as follows:
$\mathcal{L}_{ce}=-\frac{1}{B}\sum^{B}_{j=1}\mathbf{Y}^{m}_{j}\mathrm{log}(\mathbf{P}^{m}(\mathbf{k}^{m}_{j},\Theta^{m}_{p})),$ (13)
where $B$ represents the batch size, and $\mathbf{Y}^{m}_{j}=\{\mathbf{Y}^{m}_{j,tran},\mathbf{Y}^{m}_{j,rot},\mathbf{Y}^{m}_{j,grip},\mathbf{Y}^{m}_{j,col}\}$ is the ground truth for predicting the next keyframe. In light of this, NBCagent can continually learn skill-wise knowledge from current dataset $\mathcal{T}^{m}$ and memory buff $\mathcal{M}$. However, due to the limited amount of data available from old skills in memory buff $\mathcal{M}$, the data imbalance occurs between novel and old skills, further resulting in overfitting to learn novel skill-wise knowledge and forgetting skill-shared knowledge on old skills.
To address the aforementioned problems, we take an attempt to employ knowledge distillation in NERL problem. Specifically, we initialize a teacher model with the perceiver model $\Theta^{m-1}_{p}$ from the last task to extract the soft label: $\hat{\mathbf{Y}}^{m}_{j}=\mathbf{P}^{m}(\mathbf{k}^{m}_{j},\Theta^{m-1}_{p}),\hat{\mathbf{Y}}^{m}_{j}=\{\hat{\mathbf{Y}}^{m}_{j,tran},\hat{\mathbf{Y}}^{m}_{j,rot},\hat{\mathbf{Y}}^{m}_{j,grip},\hat{\mathbf{Y}}^{m}_{j,col}\}$ and apply the Kullback-Leibler divergence to align the outputs of two agents as follows:
$\mathcal{L}_{srd}=\frac{1}{\hat{B}}\sum^{B}_{j=1}\rho(\hat{\mathbf{Y}}^{m}_{j}/\tau)\mathrm{log}(\frac{\rho(\hat{\mathbf{Y}}^{m}_{j}/\tau)}{\rho(\mathbf{P}^{m}(\mathbf{k}^{m}_{j},\Theta^{m}_{p})/\tau)})\cdot\mathbb{I}_{\mathbf{k}^{m}_{j}\notin\mathcal{T}^{m}},$ (14)
where $\hat{B}=\sum_{j=1}^{B}\mathbb{I}_{\mathbf{k}^{m}_{j}\notin\mathcal{T}^{m}}$, and $\tau$ is a temperature hyper-parameter.
In conclusion, to perform skill-specific knowledge learning, we first develop SEP to accumulate skill-specific knowledge on latent and low-rank space, thereby effectively learning novel skills. Furthermore, SSR and SRD modules are designed to transfer skill-shared knowledge from semantics and representation aspects, resulting in efficiently tackling old skill forgetting. The optimization of our NBCagent can be simplified as:
$\mathcal{L}_{total}=\mathcal{L}_{ce}+\mathcal{L}_{srr}+\lambda_{2}\mathcal{L}_{srd}.$ (15)
@@ 3.4 Skill-Shared Representation Distillation Module
4 Experiments
4.1 Implementation Details
Dataset. Following PerAct, we conduct our experiments on RLBench (James et al., 2020) and simulate in CoppelaSim (Rohmer et al., 2013). To simulate the working scenarios of NERL, we design two NERL benchmark datasets, called Kitchen and Living Room. Specifically, Kitchen is constructed by gathering 10 manipulation skills pertinent to kitchen environments, and Living Room consists of 12 manipulation skills associated with living room scenarios. Each manipulation skill includes a training set of 20 episodes and a test set of 25 episodes. Furthermore, these skills involve various variations encompassing randomly sampled attributes such as colors, sizes, counts, placements, and object categories, resulting in a total of 101 distinct variations. We provide comprehensive details of two benchmark datasets in Appendix A.
Table 2: Comparison results on Kitchen under the setting of 5-1. Red and Blue represents the highest results and runner-up.
Comparison Methods 1 2 3 4 5 6 7 8 9 10 All Imp.

### v2

@@ 3.2 Skill-Shared Semantic Rendering Module
0:  Continual skill learning tasks $\{\mathcal{T}^{m}\}_{m=1}^{M}$ with datasets $\mathcal{T}^{m}=\{\mathcal{D}^{m}\}_{i=1}^{N^{d}}$; Initialized perceiver model: $\Theta^{0}_{p}$; Initialized NeRF model: $\Theta^{0}_{n}$; Pre-trained diffusion model: $\Theta_{u}$; Pre-trained CLIP language encoder: $\mathcal{E}_{c}$; Initialized memory buff: $\mathcal{M}=\emptyset$; Iterations: $\{\mathcal{I}^{m}\}_{m=1}^{M}$.
@@
5:     Compute $\mathcal{L}_{\mathrm{SSR}}$ by SSR ($\mathrm{v}^{m},\mathrm{v}^{m-1},\mathrm{l}_{x},\Theta_{u}$) using Eq. (1);
6:     $\mathbf{S}[h,:],\mathbf{W}[h,:]\xleftarrow{}$ SEP ($\mathbf{l}_{s}$);
7:     Compute $\mathcal{L}_{\mathrm{CE}},\mathcal{L}_{\mathrm{SRD}}$ utilizing $\mathbf{S}[h,:]$, $\mathbf{W}[h,:]$, $\Theta^{m}_{p},\Theta^{m-1}_{p}$;
@@
3.3 Skill-Shared Representation Distillation Module
To address semantic overlooking on past learned skills, we develop a skill-shared representation distillation module (SRD) to align skill-shared semantic in 3D voxel representation space, as presented in Fig. 1. Specifically, given a multi-modal keyframe input $\mathbf{k}^{m}_{j}=\{\mathbf{a}^{m}_{j},\mathbf{r}^{m}_{j},\mathbf{l}^{m}\}$ from a mini-batch, we feed it into our perceiver model $\mathrm{\Theta}^{m}_{p}$ to obtain a keyframe on voxel space, denoted as voxel representation $\mathbf{v}^{m}_{r,j}$. The next keyframe prediction can then be computed utilizing a Q-function head as $\mathbf{Q}^{m}(\mathbf{v}^{m}_{r,j})=\{\mathbf{P}^{m}_{j,tran},\mathbf{P}^{m}_{j,rot},\mathbf{P}^{m}_{j,grip},\mathbf{P}^{m}_{j,col}\}$, where $\mathbf{P}^{m}_{j,tran},\mathbf{P}^{m}_{j,rot},\mathbf{P}^{m}_{j,grip},\mathbf{P}^{m}_{j,col}$ denote prediction on discretized translation, rotation, gripper open state and collision avoidance. To supervise NBAgent to learn skill-wise knowledge, we follow [39], and introduce a cross-entropy loss as :
$\mathcal{L}_{\mathrm{CE}}=-\frac{1}{B}\sum^{B}_{j=1}\mathbf{Y}^{m}_{j}\mathrm{log}(\mathbf{Q}^{m}(\mathbf{v}^{m}_{r,j}),~{}~{}\mathbf{v}^{m}_{r,j}=\Theta_{p}^{m}(k^{m}_{j}),$ (6)
where $B$ represents the batch size, and $\mathbf{Y}^{m}_{j}=\{\mathbf{Y}^{m}_{j,tran},\mathbf{Y}^{m}_{j,rot},\mathbf{Y}^{m}_{j,grip},\mathbf{Y}^{m}_{j,col}\}$ is the ground truth for predicting the next keyframe. In light of this, NBAgent can continually learn skill-wise knowledge from current dataset $\mathcal{T}^{m}$ and memory buff $\mathcal{M}$. However, due to the limited amount of data available from old skills in memory buff $\mathcal{M}$, the skill-shared semantic drift on 3D voxel representation occurs between novel and old skills. It further results in forgetting skill-shared knowledge on old skills.
To address the aforementioned problems, we take an attempt to employ knowledge distillation to align voxel representation between old and new model in NBRL problem. Specifically, we initialize a teacher model with the perceiver model $\Theta^{m-1}_{p}$ from the last task to extract the soft label: $\hat{\mathbf{Y}}^{m}_{j}=\mathbf{Q}^{m-1}(\mathbf{v}^{m-1}_{r,j}),\hat{\mathbf{Y}}^{m}_{j}=\{\hat{\mathbf{Y}}^{m}_{j,tran},\hat{\mathbf{Y}}^{m}_{j,rot},\hat{\mathbf{Y}}^{m}_{j,grip},\hat{\mathbf{Y}}^{m}_{j,col}\}$ and apply the Kullback-Leibler divergence to align the outputs of two agents as follows:
$\mathcal{L}_{\mathrm{SRD}}=\frac{1}{\hat{B}}\sum^{B}_{j=1}\rho(\hat{\mathbf{Y}}^{m}_{j}/\tau)\mathrm{log}(\frac{\rho(\hat{\mathbf{Y}}^{m}_{j}/\tau)}{\rho(\mathbf{Q}^{m}(\mathbf{v}^{m}_{r,j}/\tau)})\cdot\mathbb{I}_{\mathbf{k}^{m}_{j}\notin\mathcal{T}^{m}},$ (7)
where $\hat{B}=\sum_{j=1}^{B}\mathbb{I}_{\mathbf{k}^{m}_{j}\notin\mathcal{T}^{m}}$, and $\tau$ is a hyper-parameter represents distillation temperature.
3.4 Skill-Specific Evolving Planner
To learn category-wise knowledge over the above 3D scene representation, existing continual learning methods [33, 12] assume that the knowledge acquired from novel tasks and that from previous tasks are mutually independent. Differently, we consider decoupling the knowledge as the skill-shared knowledge and skill-specific knowledge, when learning skill-wise knowledge. For instance, an agent aims to “stack the wine bottle” after learning how to “open the wine bottle”. It does not need to relearn the skill-shared knowledge of 3D scene understanding and object recognition; instead, the agent is expected to focus on acquiring the skill-specific knowledge related to the operation sequence of stacking and reducing forgetting on skill-shared knowledge. Considering this motivation, we design a skill-specific evolving planner (SEP) to perform knowledge decoupling, enabling effectively continual learning of novel manipulation skills. Specifically, inspired by [47], we first develop an adaptive language semantic bank to retrieve the skill-specific language semantic embeddings. As for these skill-specific embeddings, SEP can effectively guide the perceiver model $\Theta^{m}_{p}$ to encode multi-modal input from skill-specific latent and low-rank space. It results in learning the novel knowledge through the above skill-shared backbone and a light skill-specific subnetwork.
When observing a novel skill, we utilize a language encoder $\mathcal{E}_{c}$ from CLIP [32] to encode the input language instruction $\mathbf{l}$. Then a skill-specific language semantic embedding $\mathbf{l}_{s}\in\mathbb{R}^{D^{s}}$ can be obtained as: $\mathbf{l}_{s}=\mathcal{E}_{c}(\mathbf{l})$, where $D^{s}$ represents the dimension of $\mathbf{l}_{s}$. To build an adaptive language semantic bank $\mathcal{B}$, we then compensate the semantic embedding $\mathbf{l}_{s}$ via an exponential moving average strategy:
$\mathcal{B}[h,:]=(1-\mathcal{C}_{max})\mathcal{B}[h,:]+\mathcal{C}_{max}\mathbf{l}_{s},$ (8)
where $\mathcal{B}\in\mathbb{R}^{N^{b}\times D^{s}}$ is initialized by $N^{b}$ zero vectors. $\mathcal{C}\in\mathbb{R}^{N^{b}}$ is the cosine similarity matrix between the sentence information $\mathbf{l}_{s}$ and each vector in $\mathcal{B}$. if $\mathcal{C}_{max}>\delta$, we set $h=\mathrm{argmax}(\mathcal{C})$; otherwise, we set $h=\mathrm{nonzero}(\mathcal{B})+1$ and $\mathcal{C}_{max}=1$, where $\delta$ is a fixed threshold, and $\mathrm{nonzero}(\cdot)$ is an operation employed to calculate the number of nonzero vectors in $\mathcal{B}$. In light of this, each skill corresponds to a skill-wise code $h$.
Algorithm 2 Pipeline of Our SEP.
0:  language embedding $\mathbf{l}_{s}$;
0:  $\mathbf{S}[h,:],\mathbf{W}[h,:]$;
2:  if $\mathcal{C}_{max}>\delta$ then
3:     $h\xleftarrow{}\mathrm{argmax}(\mathcal{C});$
4:     Update $\mathcal{B}$ by Eq. (8);
5:     Return: $\mathbf{S}[h,:],\mathbf{W}[h,:]$;
6:  else
7:     $h\xleftarrow{}\mathrm{nonzero}(\mathcal{B}+1);$
8:     Expand $\mathcal{B}$ by Eq. (8);
9:     Randomly initialize $\mathbf{S}[h,:],\mathbf{W}[h,:]$;
10:     Return: $\mathbf{S}[h,:],\mathbf{W}[h,:]$;
11:  end if
To better learn skill-specific knowledge, we utilize the skill-wise code $h$ to guide the network to perform skill-specific network training. Although existing multi-modal behaviour-cloning methods [39, 53] encode multi-modal input from a skill-shared latent space, we here consider developing a dynamic skill-specific latent space $\mathbf{S}\in\mathbb{R}^{N^{s}\times N^{l}\times D}$ to facilitate skill-specific knowledge learning, where $N^{s}$, $N^{l}$, $D$ denote the number of learned skills, the learnable latent vector quantity and the vector dimension. By following [39], we encode these latent vectors in $\mathbf{S}$ with the multi-modal input to obtain the cross-attention features $\mathbf{F}_{c}\in\mathbb{R}^{(N^{p}+N^{e})\times D}$ as follows:
$\!\!\mathbf{F}_{c}=\rho(\frac{\mathrm{Cat}(\mathbf{p},\mathbf{e})\mathbf{W}_{q}(\mathbf{S}[h,:]\mathbf{W}_{k})^{\top}}{\sqrt{d}})(\mathbf{S}[h,:]\mathbf{W}_{v}),$ (9)
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{D\times D}$ are linear projection layers, and $\mathrm{Cat}(\cdot)$ denotes the concatenation operation. $\rho$ indices the softmax function and $d$ is a scaling factor. $\mathbf{p}\in\mathbb{R}^{N^{p}\times D}$ is obtained by concatenating the current state of agent and the input voxelized observation. $\mathbf{e}\in\mathbb{R}^{N^{e}\times D}$ is the encoded embeddings of the input language instruction. To this end, relying on the skill-wise code $h$, NBAgent can continually encode the novel multi-modal input from a skill-specific latent space, which is beneficial to learn skill-specific knowledge from latent space.
Considering the limitation in representing skill-specific knowledge from latent space, we further explore to learn skill-specific knowledge from low-rank space. Specifically, we introduce a low-rank adaptation layer (LoRA) [19] that can learn skill-specific knowledge in an efficient manner. For $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v},\mathbf{W}_{o}$ in each attention block, we design a set of skill-specific LoRA layers to perform skill-specific forward and obtain the final output feature $\mathbf{F}_{x}$ as follows:
$\mathbf{F}_{x}=\mathbf{X}\mathbf{W}+\mathbf{X}\mathbf{W}_{r}[h,:]=\mathbf{X}\mathbf{W}+\mathbf{X}\mathbf{W}_{a}[h,:]\mathbf{W}_{b}[h,:],$ (10)
where $\mathbf{W}\in\mathbb{R}^{D\times D}$ represents one of $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v},\mathbf{W}_{o}$ and $\mathbf{X}\in\mathbb{R}^{(N^{p}+N^{e})\times D}$ denotes the input feature for these projection layers. $\mathbf{W}_{r}[h,:]=\mathbf{W}_{a}[h,:]\mathbf{W}_{b}[h,:]$ is a low-rank decomposition, where $\mathbf{W}_{a}\in\mathbb{R}^{D\times N^{r}}$ is initialized in a random Gaussian manner, $\mathbf{W}_{b}\in\mathbb{R}^{N^{r}\times D}$ is initialized by zero and $N^{r}$ is a hyper-parameter controlling the size of LoRA layers. On the one hand, obviously, $\mathbf{W}$ is shared among all skills and expected to learn skill-shared knowledge. On the other hand, each $\mathbf{W}_{r}$ is executed to learn the corresponding skills and perform skill-specific forward with the guidance of skill-wise code $h$, resulting in continually embedding skill-specific knowledge to our NBAgent from low-rank space. Specifically, we summary the process of our SEP in Algorithm 2.
@@ 3.4 Skill-Specific Evolving Planner
Figure 2: Visualization of prediction results on various manipulation skills between Ours, Ours-w/oSRD, Ours-w/oSEP&SRD, ER and Fine-Tuning.
In conclusion, both SSR and SRD modules above are designed to capture the 3D geometric and semantic information related to the skill, and further transfer skill-shared semantic amongst diverse manipulation skills to tackle past skill forgetting. To perform skill-specific knowledge learning, we develop the SEP to accumulate skill-specific knowledge in latent and low-rank space, thereby effectively learning novel skills. Finally, the objective function of our NBAgent can be simplified as:
$\mathcal{L}_{\mathrm{Total}}=\mathcal{L}_{\mathrm{CE}}+\mathcal{L}_{\mathrm{SSR}}+\lambda_{2}\mathcal{L}_{\mathrm{SRD}}.$ (11)
4 Experiments
4.1 Implementation Details
Dataset. By following PerAct [39], we conduct our experiments on RLBench [22] and simulate robot learning in CoppelaSim [34]. To simulate the working scenarios of NBRL, we design two NBRL benchmark datasets, called Kitchen and Living Room. Specifically, Kitchen is constructed by gathering 10 manipulation skills pertinent to kitchen environments, and Living Room consists of 12 manipulation skills associated with living room scenarios. Each manipulation skill includes a training set of 20 episodes and a test set of 25 episodes. Furthermore, these skills involve various variations encompassing randomly sampled attributes such as colors, sizes, counts, placements, and object categories, resulting in a total of 101 distinct variations. In this paper, we use fixed skill ID 1-10 and 1-12 to indicate each skill in different scenarios. The correspondence between each skill ID and the details of skills are provided in Appendix A.1.
Baselines. We conduct the comprehensive evaluation between our NBAgent and the following four methods: PerAct [39], GNFactor [53], Fine-Tuning and ER [7]. PerAct and GNFactor joint train all manipulation skills within one-stage training, for two datasets respectively, referred to as the upperbound. Fine-Tuning achieves continual learning by fine-tuning all parameters in perceiver model on novel skills. ER randomly stores old skills data to memory buff and replay them when detecting novel skills.
@@ 4.1 Implementation Details
Table 1: Comparisons of success rate (%) on Kitchen and Living Room. Red and Blue represents the highest results and runner-up. The number of steps equals the sum of quantity of base skill learning task and continual skill learning tasks.
@@
Training Details. In NBRL, following continual learning setting [12], we assume that the agent undergoes initial learning through a set of manipulation skills, referred to as the base skill learning task. Subsequently, skills are added incrementally in steps, characterized as continual skill learning tasks. To valuate the robustness of NBAgent in addressing NBRL problem, we set up four combinations of base skills and new skills: 5-5, 5-1, 6-3, and 6-2, which are highlighted in bold in Tab. 1. For instance, ’5-1’ indicates that the agent is initially trained on a base skill learning task comprising the first five manipulation skills in the skill ID. Subsequently, the agent undergoes five continual skill learning tasks, each involving a new skill, ultimately learning a total of 10 skills in the Kitchen dataset. In addition, the LAMB [51] optimizer is applied for all methods with a initial learning rate of $5.0\times 10^{-4}$ and a batch size of 2. We utilize 100K training iterations for PerAct and GNFactor, 80K for base task training, 20K for incremental task training. We randomly store a fixed 4 episodes of each old skill in $\mathcal{M}$ for ER and Ours. Additional training details are available in Appendix A.2.
Evaluation Metric. Following behaviour-cloning methods [53, 17], we adopt the success score (%) $\mathcal{I}^{m}_{i}$ as the basic indicator for evaluation, where $\mathcal{I}^{m}_{i}$ represents the success score of $i$-th skill tested at $m$-th continual skill learning task. Specifically, after learning the final skill learning task, we compute the mean success score of the skills learned at the base skill learning task (Base) $\mathcal{I}^{M}_{B}$, continual skill learning task (Novel) $\mathcal{I}^{M}_{N}$ and all skill learning tasks (All) $\mathcal{I}^{M}_{A}$. These metrics respectively reflect the robustness of old skill forgetting , the capacity of novel skill learning, as well as its overall performance. Additionally, we introduce a Avg. metric $G$ and For. metric $F$ to measure average performance and skill-wise forgetting rate over the whole NBRL process, where $G=\frac{1}{M}\sum_{m=1}^{M}\mathcal{I}^{m}_{A}$ and $F=\frac{1}{N^{m}}\sum^{N^{m}}_{i=1}\underset{m\in\{1,...,M-1\}}{\mathrm{max}}(\mathcal{I}^{m}_{i}-\mathcal{I}^{M}_{i})$.
@@
4.2 Comparison Performance
We present comparison results between NBAgent and other methods on Kitchen and Living Room datasets in Tabs. 1, 3, 4 and Fig. 2. As shown in Tab. 1, NBAgent significantly outperforms compared methods by $2.5\%\sim 41.4\%$ in terms of Avg metric on four NBRL settings. This indicates that NBAgent effectively addressing old skill forgetting through skill-specific and skill-shared knowledge learning. Furthermore, we presented the skill-wise comparison results in Tabs. 3 and 4. Our model exhibits the highest success rate across all manipulation skills in terms of All. metric, improving by $4.4\%\sim 36.3\%$ and $6.2\%\sim 25.6\%$ on two datasets respectively, compared with Fine-Tuning and ER. This further demonstrates the effectiveness of our model in addressing the NBRL problem. Additionally, our NBAgent is lower than ER by $6.7\%$ in terms of For metrics on Living Room 6-2 setting, which suggests the robust and significant performance of NBAgent over the whole NBRL process. Surprisingly, as shown in Tabs. 3 and 4, our NBAgent performs better than joint training methods, specifically PerAct and GNFactor, providing additional evidence to support the efficacy of our model. The additional comparison results under other settings can be found in Appendix A.3.
Buffer size 6-3 $(2~{}\mathrm{steps})$ 6-2 $(4~{}\mathrm{steps})$
1-6 7-12 All Avg. For. 1-6 7-12 All Avg. For.
$|\mathcal{M}|=2$ 42.0 41.3 41.7 44.9 4.4 32.0 30.7 31.3 37.2 18.8
$|\mathcal{M}|=4$ 44.9 42.2 43.6 47.6 7.7 45.8 35.3 40.6 43.5 10.9
$|\mathcal{M}|=6$ 50.0 36.7 43.3 45.0 2.2 50.0 34.7 42.3 45.4 8.0
4.3 Ablation Studies
Ablation on skill-shared attribute. Compared to Ours, the scores of Ours-w/oSRD on old skills are dropped by $3.8\%\sim 12.5\%$ (i.e., 1-5 and 1-6 in Tab. 1). This indicates that SRD can effectively learn skill-shared knowledge to tackle semantic overlooking on old skills, To evaluate the effectiveness of SSR module, we visualize the rendering results in Fig. 3. It suggests that SSR module can efficiently transfer skill-shared semantics of 3D scenes, thereby achieving an improvement about $0.5\%\sim 3.1\%$ in terms of Avg compared Ours-w/oSEP&SRD with ER in Tab. 1. The above mentioned demonstrates that NBAgent can significantly tackle old skill forgetting by preserving skill-shard semantics of 3D scenes and voxel representation.
Ablation on skill-specific attribute. To evaluate effectiveness of our NBAgent on learning skill-specific knowledge, we eliminate our proposed SEP module and present results in Tabs. 1, 3 and 4. Ours-w/oSRD outperforms Ours-w/oSEP&SRD on novel skills by $1.3\%\sim 11.4\%$, (i.e., 6-10 and 7-12 in Tab. 1), which demonstrates that SEP benefits our model in learning novel skills by performing skill-specific knowledge learning.
Ablation on memory size.We also explore the impact of various sizes of memory buffer $\mathcal{M}$. As shown in Tab. 2, with $|\mathcal{M}|=6$, the forgetting rate of our model notably reduced by $1.1\%\sim 10.8\%$ in comparison to $|\mathcal{M}|=4$ and $2$, respectively. This indicates that increasing the memory size significantly addresses catastrophic forgetting but also incurs a larger memory load.
Table 3: Comparison results on Kitchen under the setting of 5-1.
Skill ID 1 2 3 4 5 6 7 8 9 10 All Imp.